A 4.0 cm² in the x-y plane sits in a uniform electric field E = (2.0 i + 3.0 j +50 k) N/C. Find the electric flux through the square.



1
Expert's answer
2021-05-30T13:28:54-0400

"\Phi=\int\vec E\cdot\vec n\cdot dA"


"\vec n=\vec k"


"\vec E\cdot\vec n=(2\vec i+3\vec j+50\vec k)\cdot\vec k=50\ (N/C)"


"\Phi=50\cdot0.0004=0.02\ (N\cdot m^2/C)" . Answer